OpenStudy (anonymous):

Find the inclination θ (in radians AND degrees) of the line: 4x-2y-4=0

OpenStudy (anonymous):

you got the slope?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

put it like y=mx+c

OpenStudy (anonymous):

-2y=-4x+4 ---->y=2x-2

OpenStudy (anonymous):

slope=2 then

OpenStudy (anonymous):

ok then what?

OpenStudy (anonymous):

slope=tan theta...theta=tan inverse(2)

OpenStudy (anonymous):

then since \(\tan(\theta)=2\) you can find \(\theta\) via \(\theta =\tan^{-1}(2)\) with a calculator

OpenStudy (anonymous):

63.43 degrees